# Multilingual State Space Models for Structured Question Answering in Indic Languages

### v2

@@ 1 Introduction
In recent years, the field of Natural Language Processing (NLP) has witnessed significant advancements with the development of models capable of handling complex linguistic structures and long-range dependencies. State Space Models (SSMs) Goel et al. (2022) have emerged as a promising alternative to traditional architectures like Transformers  Vaswani et al. (2017), particularly in tasks requiring efficient long-sequence modeling. Models such as Mamba  Gu and Dao (2024) and its variants have demonstrated the ability to manage long-context language modeling with constant memory usage, addressing some limitations inherent in Transformer-based models  Nguyen et al. (2024)  Bourdois (2022). Despite these advancements, the development of question-answering (QA) systems for low-resource languages, particularly Indic languages, remains a challenging endeavor. India has a population of 1.4 billion and is home to 122 languages and 270 mother tongues. Indian languages fall in the minority as far as NLP models are concerned. One of the Indian languages—Hindi—is spoken by 577.8 million people worldwide  Sabane et al. (2024). Likewise, Marathi, another Indian language, is ranked 11th on the list of most spoken languages. However, the scarcity of annotated datasets and neural models has hindered progress in building effective QA systems for these languages. Recent efforts have been made to address this gap by introducing large-scale QA datasets for Indic languages, such as the Indic QA dataset  Singh et al. (2024), which provides a substantial resource for developing systems tailored to the linguistic nuances of Indic languages. This paper explores the various SSM models to build efficient and contextually aware QA systems for Indic languages. SSMs are particularly suited for this task due to their ability to model both long-term and short-term dependencies in sequential data, making them well-equipped to handle the rich morphology, complex syntax, and contextual intricacies characteristic of Indian languages. The main contributions of this paper are:
@@ 2.1 Advances in State Space Models
The  eq. 1  Gu et al. (2022) describes a framework where the state vector $\mathbf{x}_{t}$ at time $t$ represents latent variables encapsulating the system’s underlying dynamics. The state evolution from $t$ to $t+1$ is governed by the state transition matrix $\mathbf{A}$, while the input matrix $\mathbf{B}$ determines the effect of the control input $\mathbf{u}_{t}$ on the state transitions. The control input $\mathbf{u}_{t}$ serves as an external influence, allowing modifications to the state dynamics at time $t$. Additionally, process noise $\mathbf{w}_{t}$ is modeled as a Gaussian random variable with zero mean ($\mathcal{N}(0,\mathbf{Q})$) and covariance matrix $\mathbf{Q}$, capturing uncertainties and randomness in the state transitions. The covariance matrix $\mathbf{Q}$ quantifies the variability of the process noise $\mathbf{w}_{t}$, ensuring the model accounts for inherent stochasticity.
@@ 3 Methodology
Our methodology focuses on developing a question-answering framework tailored for Indic languages by leveraging the capabilities of SSMs. Specifically, we experimented with various models, including Mamba, Mamba-2,Falcon Mamba, Jamba, Zamba, Samba, and Hymba, to address the linguistic diversity and complexity of Indic languages. Each model was evaluated for its ability to process questions, answers, and context efficiently while preserving the grammatical and semantic nuances of the respective languages. The overall framework for this system is illustrated in  fig. 1.
@@ 3.1 Dataset Preprocessing
Dataset preprocessing is a critical step in preparing the data for the SSM-based QA model. The preprocessing pipeline involves tokenization, vocabulary construction, encoding as mentioned in  section 3.1.1, and data transformation as mentioned in  section 3.1.2.
@@ 3.1.1 Tokenization
For our data preprocessing pipeline, we employed IndicNLPtokenizer  Kunchukuttan (2020), which is specifically designed to address the unique grammatical structures and linguistic characteristics of Indic languages. This tokenizer is particularly effective for handling the complex and diverse scripts of languages such as Hindi and Marathi. By maintaining consistent tokenization across the context (context about the question), question, and answer in our dataset, it enabled uniformity and alignment throughout the preprocessing pipeline. The tokenizer was trained on our corpus to capture language-specific patterns and generate a custom vocabulary tailored to Indic languages. This vocabulary included frequently occurring tokens and special tokens such as padding (<pad>), unknown tokens (<unk>), start-of-sentence (<sos>), and end-of-sentence (eos) markers, which are essential for enabling sequence alignment, handling missing or unknown words, and defining sentence boundaries during both training and inference. These functionalities ensure that the model can process variable-length inputs, manage incomplete data, and correctly interpret input sequences. After constructing the vocabulary, we performed encoding, where the starting position (answer_start) of the answer, representing the character index of the answer within the context, was mapped to its corresponding tokenized representation. This process generated the start (token_start) and end (token_end) token positions, which define the answer span within the tokenized sequence. These fields are essential for enabling the model to accurately learn and predict answer spans, ensuring precise alignment between the original text and its tokenized format for effective training and evaluation. The overall pipeline is illustrated in  fig. 2.
@@ 3.2 Training
After data preprocessing, the next step was training. Initially, we evaluated the model’s performance without any fine-tuning, using only tokenization mentioned in  section 3.1.1. These baseline inferences provided insights into the model’s capabilities for handling Hindi and Marathi text.
Subsequently, we fine-tuned the SSM model for question answering in Indic languages, we designed a training pipeline to enhance the performance for question answering in Indic languages. The fine-tuning process employed a resource-efficient strategy using low-rank adaptation (LoRA) Hu et al. (2021). LoRA employs a strategy of updating only the parameters in the projection layers of attention modules and the embeddings layer, leaving the rest of the model unchanged. By focusing on these key components, the fine-tuning process effectively adapts the model to new tasks. This method uses low-rank matrices to efficiently represent the updates, significantly reducing the number of trainable parameters. Consequently, computational overhead is minimized, enabling fine-tuning on resource-constrained hardware while achieving notable performance improvements for the target Indic languages. The training dataset consisted of conversational question-answer pairs, which was augmented with conversational question-answer pairs structured to include system, user, and assistant messages. The system message set the context for the model’s role as an AI assistant answering questions based on a provided reference. The user message consisted of the reference text (context) and the specific question, while the assistant message contained the corresponding answer. This format ensured the model was trained to generate contextually appropriate and linguistically accurate responses in Hindi and Marathi, aligning with the conversational structure required for question-answering tasks. The implementation details of the fine-tuning process, including hyperparameters and training configurations, are discussed later in  section 4.2
@@ 3.4 Prompting Strategies for Indic Language based SSMs
We provide examples of prompts with both positive and negative samples in Appendix  fig. 8, showcasing their use in few-shot prompting, while examples for one-shot prompting are illustrated in  fig. 3. Our Question-Answering pipeline is based on the few-shot prompting framework  Reynolds and McDonell (2021). This approach demonstrates how leveraging paradigms from different prompting techniques leads to improved results during fine-tuning, as evidenced in the subsequent  section 4.
@@ 4.1 Dataset
We utilized the IndicQA dataset, which was constructed by aggregating multiple open-source datasets to develop a SQuAD-style resource for Hindi and Marathi. This dataset includes a significant component of approximately 28,000 samples in these languages, making it one of the largest publicly available datasets in Hindi and Marathi  Sabane et al. (2024).The datasets are summarized in  table 2.
@@ 4.2 Implementation Details
For Mamba and Mamba-2, the fine-tuning configuration included a learning rate of $2\times 10^{-4}$, a batch size of 4 (with gradient accumulation for an effective batch size of 32), 3 epochs, 100 warmup steps, and mixed precision (FP16). The Falcon Mamba model was fine-tuned with start and end predictors, cross-entropy loss, and the Adam optimizer, using a learning rate of $1\times 10^{-4}$, a batch size of 4, and 10 epochs. The Jamba model employed a LoRA configuration with an increased rank of 16 to enhance capacity and processed sequences with a context window of up to 256K tokens. The Zamba model was optimized for long sequences, using a maximum sequence length of 4096 tokens. The Samba model focused on computational efficiency, handling sequences up to 256K tokens with a batch size of 8 (effective batch size of 64 with gradient accumulation), excelling in modeling complex conversational patterns. Finally, the Hymba model incorporated learnable meta tokens and partial sliding window attention, trained with a batch size of 4 and a learning rate of $3\times 10^{-4}$. Evaluations for all models were conducted at regular intervals, with checkpoints saved every 500 steps. At the end of training, the best-performing model was selected. The validation graphs for all models are provided in  fig. 4.
@@ 4.3 Results
This section evaluates model performance on Hindi and Marathi datasets, comparing results before and after fine-tuning. Additionally,  fig. 8 includes sample model responses to highlight qualitative differences in predictions.
@@ 4.3.1 Pre-Fine-Tuning Performance: Initial Challenge
Table 3 presents model performance before fine-tuning, showing variations across different architectures. Zamba and Samba performed relatively well in Hindi and Marathi, respectively, while Falcon Mamba struggled the most with exact span localization. Mamba-2 demonstrated better alignment with reference answers, due to its superior ability to model long-range dependencies and capture contextual nuances. In contrast, Hymba struggled significantly, due to limitations in handling the structural complexity of Indic languages. Models that performed better exhibited stronger token-level alignment and more effective handling of variations, while weaker models failed to generalize effectively, particularly in Marathi, where linguistic complexity posed additional challenges. Another key observation is that models generally performed better on Hindi than Marathi, which can be attributed to differences in dataset size, linguistic complexity, and pre-training exposure. Hindi, being more widely studied, benefits from larger corpora and more refined tokenization strategies, whereas Marathi’s syntactic variations pose additional challenges. Models that struggled, such as Falcon Mamba and Hymba, lacked the ability to effectively process inflectional diversity and complex word formations, leading to lower performance. This highlights the necessity of incorporating language-specific optimizations and better pre-training strategies to improve generalization in underrepresented Indic languages.
@@ 4.3.2 Post-Fine-Tuning Performance: Improvements & Insights
Additionally, models that initially performed poorly, such as Falcon Mamba and Hymba, saw only marginal improvements, suggesting fundamental architectural limitations that fine-tuning alone could not overcome. Beyond numerical improvements, fine-tuning also contributed to better lexical and semantic alignment. Models demonstrated enhanced fluency in response generation, with improvements in phrase structure and word selection, reducing inconsistencies seen in non-fine-tuned outputs. However, challenges persisted for certain models, particularly in maintaining coherence when processing long-form text in Marathi. These findings emphasize the necessity of fine-tuning pre-trained models on domain-specific datasets to optimize their performance for low-resource languages. Models results after fine-tuning is provided in  table 4 , illustrating the extent of improvement across different architectures.
@@ Appendix A Dataset Statistics
The scatter plot in  fig. 5 illustrates the relationship between question length and answer length across the Hindi (blue) and Marathi (red) datasets. The visualization highlights a clear positive correlation between the lengths of questions and their corresponding answers, with longer questions often leading to longer answers. Notably, the Marathi dataset shows slightly shorter average lengths compared to the Hindi dataset, reflecting language-specific characteristics. This analysis provides insights into the structural patterns within the dataset, aiding in understanding the data distribution and informing preprocessing decisions for the SSM-based QA model.
@@
The scatter plot in  fig. 6 illustrates the relationship between the starting position of answers within the context and the total context length for the Hindi (blue) and Marathi (red) datasets. The visualization shows that as the context length increases, the starting positions of answers become more varied, indicating that longer contexts are more likely to contain answers at different positions. This trend is consistent across both datasets, although the Marathi dataset tends to have shorter contexts on average compared to Hindi. This analysis highlights the diversity in answer positions relative to context length, which is crucial for designing models capable of handling varied input structures in question-answering tasks.
@@ Appendix B Results
The results, as shown in  fig. 9, illustrate the performance metrics for the Hindi and Marathi datasets, comparing outcomes before and after fine-tuning the model. Metrics such as BLEU, ROUGE, F1, Exact Match, and BERT scores are presented for both languages, highlighting the impact of fine-tuning on model performance. The left panel represent the scores without fine-tuning, while the right panels show the improved results after fine-tuning. Across both datasets, fine-tuning consistently improves all evaluation metrics, with notable gains in Exact Match and F1 scores, indicating better alignment between predicted and ground-truth answers. The BERT scores also demonstrate significant improvements, reflecting enhanced semantic understanding. These results emphasize the importance of fine-tuning for adapting the model to the nuances of Hindi and Marathi languages in the QA task.

### v3

@@ 1 Introduction
In recent years, the field of Natural Language Processing (NLP) has witnessed significant advancements with the development of models capable of handling complex linguistic structures and long-range dependencies. State Space Models (SSMs) Goel et al. (2022) have emerged as a promising alternative to traditional architectures like Transformers  Vaswani et al. (2017), particularly in tasks requiring efficient long-sequence modeling. Models such as Mamba  Gu and Dao (2024) and its variants have demonstrated the ability to manage long-context language modeling with constant memory usage, addressing some limitations inherent in Transformer-based models  Nguyen et al. (2024)  Bourdois (2022). Despite these advancements, the development of question-answering (QA) systems for low-resource languages, particularly Indic languages, remains a challenging endeavor. India has a population of 1.4 billion and is home to 122 languages and 270 mother tongues. Indian languages fall in the minority as far as NLP models are concerned. One of the Indian languages—Hindi—is spoken by 577.8 million people worldwide  Sabane et al. (2024). Likewise, Marathi, another Indian language, is ranked 11th on the list of most spoken languages  Raja and Vats (2025). However, the scarcity of annotated datasets and neural models has hindered progress in building effective QA systems for these languages. Recent efforts have been made to address this gap by introducing large-scale QA datasets for Indic languages, such as the Indic QA dataset  Singh et al. (2024), which provides a substantial resource for developing systems tailored to the linguistic nuances of Indic languages. This paper explores the various SSM models to build efficient and contextually aware QA systems for Indic languages. SSMs are particularly suited for this task due to their ability to model both long-term and short-term dependencies in sequential data, making them well-equipped to handle the rich morphology, complex syntax, and contextual intricacies characteristic of Indian languages. The main contributions of this paper are:
@@ 2.1 Advances in State Space Models
The  eq.˜1  Gu et al. (2022) describes a framework where the state vector $\mathbf{x}_{t}$ at time $t$ represents latent variables encapsulating the system’s underlying dynamics. The state evolution from $t$ to $t+1$ is governed by the state transition matrix $\mathbf{A}$, while the input matrix $\mathbf{B}$ determines the effect of the control input $\mathbf{u}_{t}$ on the state transitions. The control input $\mathbf{u}_{t}$ serves as an external influence, allowing modifications to the state dynamics at time $t$. Additionally, process noise $\mathbf{w}_{t}$ is modeled as a Gaussian random variable with zero mean ($\mathcal{N}(0,\mathbf{Q})$) and covariance matrix $\mathbf{Q}$, capturing uncertainties and randomness in the state transitions. The covariance matrix $\mathbf{Q}$ quantifies the variability of the process noise $\mathbf{w}_{t}$, ensuring the model accounts for inherent stochasticity.
@@ 3 Methodology
Our methodology focuses on developing a question-answering framework tailored for Indic languages by leveraging the capabilities of SSMs. Specifically, we experimented with various models, including Mamba, Mamba-2,Falcon Mamba, Jamba, Zamba, Samba, and Hymba, to address the linguistic diversity and complexity of Indic languages. Each model was evaluated for its ability to process questions, answers, and context efficiently while preserving the grammatical and semantic nuances of the respective languages. The overall framework for this system is illustrated in  fig.˜1.
@@ 3.1 Dataset Preprocessing
Dataset preprocessing is a critical step in preparing the data for the SSM-based QA model. The preprocessing pipeline involves tokenization, vocabulary construction, encoding as mentioned in  section˜3.1.1, and data transformation as mentioned in  section˜3.1.2.
@@ 3.1.1 Tokenization
For our data preprocessing pipeline, we employed IndicNLPtokenizer  Kunchukuttan (2020), which is specifically designed to address the unique grammatical structures and linguistic characteristics of Indic languages. This tokenizer is particularly effective for handling the complex and diverse scripts of languages such as Hindi and Marathi. By maintaining consistent tokenization across the context (context about the question), question, and answer in our dataset, it enabled uniformity and alignment throughout the preprocessing pipeline. The tokenizer was trained on our corpus to capture language-specific patterns and generate a custom vocabulary tailored to Indic languages. This vocabulary included frequently occurring tokens and special tokens such as padding (<pad>), unknown tokens (<unk>), start-of-sentence (<sos>), and end-of-sentence (eos) markers, which are essential for enabling sequence alignment, handling missing or unknown words, and defining sentence boundaries during both training and inference. These functionalities ensure that the model can process variable-length inputs, manage incomplete data, and correctly interpret input sequences. After constructing the vocabulary, we performed encoding, where the starting position (answer_start) of the answer, representing the character index of the answer within the context, was mapped to its corresponding tokenized representation. This process generated the start (token_start) and end (token_end) token positions, which define the answer span within the tokenized sequence. These fields are essential for enabling the model to accurately learn and predict answer spans, ensuring precise alignment between the original text and its tokenized format for effective training and evaluation. The overall pipeline is illustrated in  fig.˜2.
@@ 3.2 Training
After data preprocessing, the next step was training. Initially, we evaluated the model’s performance without any fine-tuning, using only tokenization mentioned in  section˜3.1.1. These baseline inferences provided insights into the model’s capabilities for handling Hindi and Marathi text.
Subsequently, we fine-tuned the SSM model for question answering in Indic languages, we designed a training pipeline to enhance the performance for question answering in Indic languages. The fine-tuning process employed a resource-efficient strategy using low-rank adaptation (LoRA) Hu et al. (2021). LoRA employs a strategy of updating only the parameters in the projection layers of attention modules and the embeddings layer, leaving the rest of the model unchanged. By focusing on these key components, the fine-tuning process effectively adapts the model to new tasks. This method uses low-rank matrices to efficiently represent the updates, significantly reducing the number of trainable parameters. Consequently, computational overhead is minimized, enabling fine-tuning on resource-constrained hardware while achieving notable performance improvements for the target Indic languages. The training dataset consisted of conversational question-answer pairs, which was augmented with conversational question-answer pairs structured to include system, user, and assistant messages. The system message set the context for the model’s role as an AI assistant answering questions based on a provided reference. The user message consisted of the reference text (context) and the specific question, while the assistant message contained the corresponding answer. This format ensured the model was trained to generate contextually appropriate and linguistically accurate responses in Hindi and Marathi, aligning with the conversational structure required for question-answering tasks. The implementation details of the fine-tuning process, including hyperparameters and training configurations, are discussed later in  section˜4.2
@@ 3.4 Prompting Strategies for Indic Language based SSMs
We provide examples of prompts with both positive and negative samples in Appendix  fig.˜8, showcasing their use in few-shot prompting, while examples for one-shot prompting are illustrated in  fig.˜3. Our Question-Answering pipeline is based on the few-shot prompting framework  Reynolds and McDonell (2021). This approach demonstrates how leveraging paradigms from different prompting techniques leads to improved results during fine-tuning, as evidenced in the subsequent  section˜4.
@@ 4.1 Dataset
We utilized the IndicQA dataset, which was constructed by aggregating multiple open-source datasets to develop a SQuAD-style resource for Hindi and Marathi. This dataset includes a significant component of approximately 28,000 samples in these languages, making it one of the largest publicly available datasets in Hindi and Marathi  Sabane et al. (2024).The datasets are summarized in  table˜2.
@@ 4.2 Implementation Details
For Mamba and Mamba-2, the fine-tuning configuration included a learning rate of $2\times 10^{-4}$, a batch size of 4 (with gradient accumulation for an effective batch size of 32), 3 epochs, 100 warmup steps, and mixed precision (FP16). The Falcon Mamba model was fine-tuned with start and end predictors, cross-entropy loss, and the Adam optimizer, using a learning rate of $1\times 10^{-4}$, a batch size of 4, and 10 epochs. The Jamba model employed a LoRA configuration with an increased rank of 16 to enhance capacity and processed sequences with a context window of up to 256K tokens. The Zamba model was optimized for long sequences, using a maximum sequence length of 4096 tokens. The Samba model focused on computational efficiency, handling sequences up to 256K tokens with a batch size of 8 (effective batch size of 64 with gradient accumulation), excelling in modeling complex conversational patterns. Finally, the Hymba model incorporated learnable meta tokens and partial sliding window attention, trained with a batch size of 4 and a learning rate of $3\times 10^{-4}$. Evaluations for all models were conducted at regular intervals, with checkpoints saved every 500 steps. At the end of training, the best-performing model was selected. The validation graphs for all models are provided in  fig.˜4.
@@ 4.3 Results
This section evaluates model performance on Hindi and Marathi datasets, comparing results before and after fine-tuning. Additionally,  fig.˜8 includes sample model responses to highlight qualitative differences in predictions.
@@ 4.3.1 Pre-Fine-Tuning Performance: Initial Challenge
Table˜3 presents model performance before fine-tuning, showing variations across different architectures. Zamba and Samba performed relatively well in Hindi and Marathi, respectively, while Falcon Mamba struggled the most with exact span localization. Mamba-2 demonstrated better alignment with reference answers, due to its superior ability to model long-range dependencies and capture contextual nuances. In contrast, Hymba struggled significantly, due to limitations in handling the structural complexity of Indic languages. Models that performed better exhibited stronger token-level alignment and more effective handling of variations, while weaker models failed to generalize effectively, particularly in Marathi, where linguistic complexity posed additional challenges. Another key observation is that models generally performed better on Hindi than Marathi, which can be attributed to differences in dataset size, linguistic complexity, and pre-training exposure. Hindi, being more widely studied, benefits from larger corpora and more refined tokenization strategies, whereas Marathi’s syntactic variations pose additional challenges. Models that struggled, such as Falcon Mamba and Hymba, lacked the ability to effectively process inflectional diversity and complex word formations, leading to lower performance. This highlights the necessity of incorporating language-specific optimizations and better pre-training strategies to improve generalization in underrepresented Indic languages.
@@ 4.3.2 Post-Fine-Tuning Performance: Improvements & Insights
Additionally, models that initially performed poorly, such as Falcon Mamba and Hymba, saw only marginal improvements, suggesting fundamental architectural limitations that fine-tuning alone could not overcome. Beyond numerical improvements, fine-tuning also contributed to better lexical and semantic alignment. Models demonstrated enhanced fluency in response generation, with improvements in phrase structure and word selection, reducing inconsistencies seen in non-fine-tuned outputs. However, challenges persisted for certain models, particularly in maintaining coherence when processing long-form text in Marathi. These findings emphasize the necessity of fine-tuning pre-trained models on domain-specific datasets to optimize their performance for low-resource languages. Models results after fine-tuning is provided in  table˜4 , illustrating the extent of improvement across different architectures.
@@ Appendix A Dataset Statistics
The scatter plot in  fig.˜5 illustrates the relationship between question length and answer length across the Hindi (blue) and Marathi (red) datasets. The visualization highlights a clear positive correlation between the lengths of questions and their corresponding answers, with longer questions often leading to longer answers. Notably, the Marathi dataset shows slightly shorter average lengths compared to the Hindi dataset, reflecting language-specific characteristics. This analysis provides insights into the structural patterns within the dataset, aiding in understanding the data distribution and informing preprocessing decisions for the SSM-based QA model.
@@
The scatter plot in  fig.˜6 illustrates the relationship between the starting position of answers within the context and the total context length for the Hindi (blue) and Marathi (red) datasets. The visualization shows that as the context length increases, the starting positions of answers become more varied, indicating that longer contexts are more likely to contain answers at different positions. This trend is consistent across both datasets, although the Marathi dataset tends to have shorter contexts on average compared to Hindi. This analysis highlights the diversity in answer positions relative to context length, which is crucial for designing models capable of handling varied input structures in question-answering tasks.
@@ Appendix B Results
The results, as shown in  fig.˜9, illustrate the performance metrics for the Hindi and Marathi datasets, comparing outcomes before and after fine-tuning the model. Metrics such as BLEU, ROUGE, F1, Exact Match, and BERT scores are presented for both languages, highlighting the impact of fine-tuning on model performance. The left panel represent the scores without fine-tuning, while the right panels show the improved results after fine-tuning. Across both datasets, fine-tuning consistently improves all evaluation metrics, with notable gains in Exact Match and F1 scores, indicating better alignment between predicted and ground-truth answers. The BERT scores also demonstrate significant improvements, reflecting enhanced semantic understanding. These results emphasize the importance of fine-tuning for adapting the model to the nuances of Hindi and Marathi languages in the QA task.